# Intersection Support for Outerplanar and Bounded Treewidth Graphs

### v1

@@ 1 Introduction
Voloshina and Feinberg introduced the notion of support as a notion of planarity of hypergraphs [voloshina1984planarity]. The notion is now well studied in hypergraph visualization [bereg2015colored, bereg2011red, brandes2010blocks, brandes2012path, buchin2011planar, havet2022overlaying, hurtado2018colored] and network design [anceaume2006semantic, baldoni2007tera, baldoni2007efficient, chand2005semantic, hosoda2012approximability, korach2003clustering, onus2011minimum]. The notion of supports and similar notions have also found applicability for geometric problems [Mustafa17, BasuRoy2018, Cohen-AddadM15, krohn2014guarding, mustafa2010improved, RR18].
@@
While the notion of support has found wide applications in different domains, we still do not have a good tool set for deciding or constructing supports even in restricted settings, many of which arise in applications. For example, Buchin et al. [buchin2011planar] showed that deciding if a hypergraph admits a 2-outerplanar [^2] A graph $G$ is 2-outerplanar if there is an embedding of $G$ in the plane such that removing the vertices of the outer face results in an outerplanar graph. support is NP-hard, and they left open the question of deciding if a hypergraph admits an outerplanar support. While we do not answer their question, we believe that the tools we develop in constructing an outerplanar support may help in eventually resolving their question.
@@ 3 Preliminaries
We now show that $\tilde{Q}$ is an intersection support for $(G,\mathcal{H,K})$. Let $K\in\mathcal{K}$ be arbitrary. We claim that $\tilde{Q}[\mathcal{H}_{K}]$ induces a connected subgraph of $\tilde{Q}$. Since $\tilde{Q}^{\prime}$ is an intersection support for $(G,\mathcal{H}^{*},\mathcal{K})$, the induced subgraph $\tilde{Q}^{\prime}[\mathcal{H}^{*}_{K}]$ of $\tilde{Q}^{\prime}$ is connected. By definition, $\mathcal{H}^{*}\subseteq\mathcal{H}$ consists of precisely the maximal elements of $(\mathcal{H},\preceq_{N})$. Therefore, for any $H\in\mathcal{H}_{K}\setminus\mathcal{H}^{*}_{K}$, $\exists H^{\prime}\in\mathcal{H}^{*}$ such that $H\subseteq H^{\prime}$ and $\{H,H^{\prime}\}$ is an edge in $\tilde{Q}$. But $H^{\prime}$ intersects $K$ since $H$ interacts $K$. It follows that $H^{\prime}\in\mathcal{H}^{*}_{K}$ and hence, the induced subgraph $\tilde{Q}[\mathcal{H}_{K}]$ of $\tilde{Q}$ is connected. This completes the proof. ∎
@@ 5.1 Cycle $axax$-free systems
To see that $(C,\mathcal{H},\mathcal{K})$ satisfies the intersection property, consider $H\in\mathcal{H}$ and $K\in\mathcal{K}$ with vertices $h_{1},k_{1},h_{2},k_{2}$ in cyclic order s.t. $h_{1},h_{2}\in H$ and $k_{1},k_{2}\in K$. Since $H$ is connected in $G$, there is a path $P$ between $h_{1}$ and $h_{2}$, all of whose vertices lie in $H$. If $H\cap K=\emptyset$, then $P$ lies in $H\setminus K$. Since $h_{1}$ and $h_{2}$ are non-consecutive on $C$, $G\setminus P$ gets separated into two components with $k_{1}$ and $k_{2}$ in distinct components. This implies there is no path between $k_{1}$ and $k_{2}$ in $G$ that lies entirely in $K$, a contradiction since $K$ is connected. ∎
@@
Let $C$ be the outer cycle in an outerplanar embedding of $G$. By Lemma 13, $(C,\mathcal{H},\mathcal{K})$ is strong $axax$-free. By Theorem 14, there is an outerplanar support $\tilde{Q}$ for $(C,\mathcal{H},\mathcal{K})$. $\tilde{Q}$ is also a support for $(G,\mathcal{H},\mathcal{K})$ since the underlying intersection hypergraph defined by $\mathcal{H}$ and $\mathcal{K}$ remains the same on $G$ and on $C$. ∎
For Theorem 6 to go through, it remains to prove Theorem 14 which we do in Section 5.2. In the below, we show that we can simplify an $axax$-free cycle system so that each vertex of $C$ is contained in at least one subgraph in $\mathcal{H}$ and one subgraph in $\mathcal{K}$.
@@ 5.2 Construction of Outerplanar Supports
In this section, we prove Theorem 14. We start with some basic terms required for the proof. Let $C=\{0,\ldots,n-1\}$ be a cycle on $n$ vertices oriented clockwise, and let $\mathcal{R}$ be a collection of arcs on $C$ such that no two arcs in $\mathcal{R}$ whose both ends are defined by vertices of $C$.
@@
Let $\mathcal{R}$ be a collection of arcs on a cycle $C$ that are strict-containment free. We construct a cycle $C(\mathcal{R})$ on $\mathcal{R}$ as follows: Starting at an arbitrary vertex on $C$, in a clockwise traversal, we construct a cycle $C(\mathcal{R})$ on $\mathcal{R}$ as follows: When we visit a vertex $v$ of $C$, for each $R\in\mathcal{R}$ s.t. $s(R)=v$, we put a vertex in $C(\mathcal{R})$. The vertices added to $C(\mathcal{R})$ w.r.t. $v$ are ordered in increasing order of $t(R)$, with ties broken arbitrarily. We call the cycle $C(\mathcal{R})$ thus constructed, the cycle in lex. cyclic order, and we say that $\mathcal{R}$ is ordered in lex. cyclic order.
@@
$C(\mathcal{H})$ is clearly an outerplanar graph. Consider a subgraph $K\in\mathcal{K}$. Since $K$ induces a single run on $C$ and the subgraphs in $\mathcal{H}$ are strict-containment free, it follows the subgraphs in $\mathcal{H}_{K}$ appear consecutively in lex. cyclic order on $C(\mathcal{H})$ and are thus induced a connected subgraph of $C(\mathcal{H})$. ∎
@@
The proof for the moreover part is similar. Indeed, let $h^{\prime}\in\mathrm{arc}[v,u]$ s.t. $h^{\prime}\in H^{\prime}\setminus H$. Also, note that $x\in H^{\prime}\setminus H$. Hence, the subgraphs $H,H^{\prime}$ form an $axax$-pattern as witnessed by the cyclic sequence $h^{\prime},u,x,v$; a contradiction. ∎
@@
Note that since $\mathcal{H}$ is strict-containment free, so are $\mathcal{H}_{R}$ and $\mathcal{H}_{L}$, and that each $H\in\mathcal{H}\setminus\{H_{1},H^{\prime}_{s}\}$ contributes a run to exactly one of $\mathcal{H}_{R}$ or $\mathcal{H}_{L}$, by construction.
@@
((b)) $(C,\mathcal{H}_{R})$
((c)) $(C,\mathcal{H}_{L})$
Figure 5: Collections $\mathcal{H}_{R}$ and $\mathcal{H}_{L}$ of derived systems. $\mathcal{K}$ subgraphs are not shown in the figure.
@@
Finally, let $K$ intersects both $\mathrm{arc}[u_{0},v_{0}]$ and $\mathrm{arc}[v_{0},u_{0}]$. Note that in this case $K$ intersects with $H_{0}$ (and hence with $H^{\prime}_{0}$ and $H^{\prime\prime}_{0}$) since $u_{0},v_{0}\in H_{0}$ and $(C,\mathcal{H},\mathcal{K})$ is strong $axax$-free. Each $H\in\mathcal{H}\setminus\{H_{0}\}$ has a representative vertex in one of $Q_{L}$ or $Q_{R}$ which are supports for the derived cycle systems with $H_{0}$ in both $Q_{L}$ and $Q_{R}$. It follows that $\mathcal{H}_{K}$ induces a connected subgraph of $Q$. Since $K$ was chosen arbitrarily, $Q$ is a support. ∎
@@ 5.3 Implementation
Hence, the VC-dimension of the set system $(C,\mathcal{H})$ is at most 4, and this implies, by the Sauer-Shelah lemma (See Chapter 10 in [Mat02LecDiscGeom]) that $|\mathcal{H}|,|\mathcal{K}|=O(n^{4})$. ∎
@@ 5.4 Primal and Dual Supports: Revisited
The existence of an intersection support also implies a primal and a dual support. However, we emphasize here there is a primal support even if the cycle system $(C,\mathcal{H})$ is just $abab$-free but not necessarily $axax$-free. This follows directly from the following result of Raman and Singh [raman2025supportsgraphsboundedgenus].
@@
For the dual setting, one may likewise hope that the $abab$-free condition is sufficient to obtain a dual support. Unfortunately, that is not the case as the following example shows: Let $G$ be the graph on $\{1,2,\ldots 6\}$ as shown in Figure 6. Let $\mathcal{H}$ be subgraphs $H_{1}=\{1,2,3\}$, $H_{2}=\{3,4,5\}$, $H_{3}=\{5,6,1\}$ and $H_{4}=\{2,4,6\}$. It can be checked that $\mathcal{H}$ is $abab$-free. The dual support for $(G,\mathcal{H})$ is $K_{4}$ which is not outerplanar.
The reason why $(G,\mathcal{H})$ does not admit a dual outerplanar support is that the induced system $(C,\mathcal{H})$ is not $axax$-free - $H_{4}$ forms an $axax$-pair with each of $H_{1},H_{2}$ and $H_{3}$.
@@ 6 Support of Bounded Treewidth
The proofs follow immediately from the fact that $\mathcal{H}$ are non-piercing and the adhesion set corresponding to any edge in $E(T)$ is a separator in $G$. ∎
@@
Since $G$ is chordal, each bag induces a complete subgraph of $G$. Hence, each $H_{x}\in\mathcal{H}_{x}$ is connected. If $x$ is not a root of $T$, let $y$ be the parent of $x$. Then, for $H_{x},H^{\prime}_{x}\in\mathcal{H}_{x}$, the subgraph $H_{x}\setminus H^{\prime}_{x}$ of $G_{x}$ is connected since $H\setminus H^{\prime}$ is connected and $A_{xy}$ is a separator in $G$. Therefore, $(G_{x},\mathcal{H}_{x})$ is a non-piercing graph system. Finally, that $(T_{x},\mathcal{B}_{x})$ is a tree decomposition, is straightforward. ∎
@@ 6.1 Primal Support
In this section, we show that a bounded treewidth non-piercing system $(G,\mathcal{H})$ with $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ admits a primal support $Q$ s.t. $\textsc{tw}(Q)\leq 2\cdot 2^{\textsc{tw}(G)}+\textsc{tw}(G)$. The proof is algorithmic and yields a polynomial time algorithm if $\textsc{tw}(G)$ is bounded, i.e., an FPT-algorithm parameterized by $\textsc{tw}(G)$.
@@
Let $(T,\mathcal{B})$ be a tree decomposition of $G$ of width $t$ where $T$ is a binary rooted tree. If $(T,\mathcal{B})$ is an easy tree decomposition, we are done by Lemma 27. Otherwise, we transform it into an easy tree decomposition of width at most $2^{t+2}+t$ by Lemma 28. We then obtain a primal support $Q$ of the same treewidth by Lemma 27. ∎
@@ 6.2 Dual Support
We now construct a dual support. For a non-piercing system $(G,\mathcal{H}_{)}$ of treewidth $t$, we construct a dual support of treewidth $O(2^{4t})$. By Lemma 5, we assume that there are no containments, i.e., there are no two subgraphs $H,H^{\prime}\in\mathcal{H}$ such that $H\subseteq H^{\prime}$. As in the primal setting, we first show how we can obtain a dual support for a simple case. A support for the general case is obtained by reducing it to the simple case. In the following, since we construct a graph on $\mathcal{H}$, we abuse notation and also use $H$ to denote the vertex in the dual support corresponding to $H\in\mathcal{H}$.
@@
We now argue that $Q^{*}$ is a dual support for $(G,\mathcal{H})$. Consider a vertex $v\in V(G)$. The algorithm above ensures that each subgraph is pushed at most once. Suppose $H^{\prime}\in\mathcal{H}_{v}$ was pushed by $H$ so that its modified copy $H^{\prime\prime}=H^{\prime}\setminus H$ does not cover $v$. Then, $H\in\mathcal{H}_{v}$. Let $e=\{a,b\}$ be the connecting edge between $H$ and $H^{\prime\prime}$ such that $a\in H$ and $b\in H^{\prime\prime}$. Since $(T,\mathcal{B})$ is a valid tree-decomposition, there is a bag $B$ containing both $a$ and $b$. Let $H_{1}$ be the unique representative of $H^{\prime\prime}$ in $\textrm{unique}(\mathcal{H}^{\prime})$. Since Algorithm $k$-SDS puts a complete graph on the subgraphs intersecting $B$, it implies $H$ and $H_{1}$ are adjacent in $Q^{\prime}$. In $Q^{*}$, we made $H^{\prime}$ adjacent to $H_{1}$. Hence, $Q^{*}$ is a dual support for $(G,\mathcal{H})$. ∎
@@
Let $(T,\mathcal{B})$ be a tree decomposition of $G$ of width $t$. If $(T,\mathcal{B})$ is $2^{4(t+1)}$-sparse, then we are done. Otherwise, by Lemma 30, we obtain a dual support $Q^{*}$ for $(G,\mathcal{H})$, of treewidth at most $2^{4(t+1)}$. ∎

### v2

# Rajiv Raman

@@ 1 Introduction
Voloshina and Feinberg introduced the notion of a support as a notion of planarity of hypergraphs [voloshina1984planarity]. The notion is now well studied in hypergraph visualization [bereg2015colored, bereg2011red, brandes2010blocks, brandes2012path, buchin2011planar, havet2022overlaying, hurtado2018colored] and network design [anceaume2006semantic, baldoni2007tera, baldoni2007efficient, chand2005semantic, hosoda2012approximability, korach2003clustering, onus2011minimum]. The notion of supports and similar notions have also found applicability for geometric problems [Mustafa17, BasuRoy2018, Cohen-AddadM15, krohn2014guarding, mustafa2010improved, RR18].
@@
While the notion of supports has found applications in different domains, we still do not have a good tool-set for deciding or constructing supports even in restricted settings, many of which arise in applications. For example, Buchin et al. [buchin2011planar] showed that deciding if a hypergraph admits a 2-outerplanar [^3] A graph $G$ is 2-outerplanar if there is an embedding of $G$ in the plane such that removing the vertices of the outer face results in an outerplanar graph. support is NP-hard, and they left open the question of deciding if a hypergraph admits an outerplanar support. While we do not answer their question, we believe that the tools we develop in constructing an outerplanar support may help in eventually resolving their question.
@@ 3 Preliminaries
We now show that $\tilde{Q}$ is an intersection support for $(G,\mathcal{H,K})$. Let $K\in\mathcal{K}$ be arbitrary. We claim that $\tilde{Q}[\mathcal{H}_{K}]$ induces a connected subgraph of $\tilde{Q}$. Since $\tilde{Q}^{\prime}$ is an intersection support for $(G,\mathcal{H}^{*},\mathcal{K})$, the induced subgraph $\tilde{Q}^{\prime}[\mathcal{H}^{*}_{K}]$ of $\tilde{Q}^{\prime}$ is connected. By definition, $\mathcal{H}^{*}\subseteq\mathcal{H}$ consists of precisely the maximal elements of $(\mathcal{H},\preceq_{N})$. Therefore, for any $H\in\mathcal{H}_{K}\setminus\mathcal{H}^{*}_{K}$, $\exists H^{\prime}\in\mathcal{H}^{*}$ such that $H\subseteq H^{\prime}$ and $\{H,H^{\prime}\}$ is an edge in $\tilde{Q}$. But $H^{\prime}$ intersects $K$ since $H$ interacts $K$. It follows that $H^{\prime}\in\mathcal{H}^{*}_{K}$ and hence, the induced subgraph $\tilde{Q}[\mathcal{H}_{K}]$ of $\tilde{Q}$ is connected. This completes the proof.
@@ 5.1 Cycle $axax$-free systems
To see that $(C,\mathcal{H},\mathcal{K})$ satisfies the intersection property, consider $H\in\mathcal{H}$ and $K\in\mathcal{K}$ with vertices $h_{1},k_{1},h_{2},k_{2}$ in cyclic order s.t. $h_{1},h_{2}\in H$ and $k_{1},k_{2}\in K$. Since $H$ is connected in $G$, there is a path $P$ between $h_{1}$ and $h_{2}$, all of whose vertices lie in $H$. If $H\cap K=\emptyset$, then $P$ lies in $H\setminus K$. Since $h_{1}$ and $h_{2}$ are non-consecutive on $C$, $G\setminus P$ gets separated into two components with $k_{1}$ and $k_{2}$ in distinct components. This implies there is no path between $k_{1}$ and $k_{2}$ in $G$ that lies entirely in $K$, a contradiction since $K$ is connected.
@@
Let $C$ be the outer cycle in an outerplanar embedding of $G$. By Lemma 13, $(C,\mathcal{H},\mathcal{K})$ is strong $axax$-free. By Theorem 14, there is an outerplanar support $\tilde{Q}$ for $(C,\mathcal{H},\mathcal{K})$. $\tilde{Q}$ is also a support for $(G,\mathcal{H},\mathcal{K})$ since the underlying intersection hypergraph defined by $\mathcal{H}$ and $\mathcal{K}$ remains the same on $G$ and on $C$.
For Theorem 6 to go through, it remains to prove Theorem 14 which we do in Section 5.2. Below, we show that we can simplify an $axax$-free cycle system so that each vertex of $C$ is contained in at least one subgraph in $\mathcal{H}$ and one subgraph in $\mathcal{K}$.
@@ 5.2 Construction of Outerplanar Supports
In this section, we prove Theorem 14. We start with some basic terminology required for the proof. Let $C=\{0,\ldots,n-1\}$ be a cycle on $n$ vertices oriented clockwise, and let $\mathcal{R}$ be a collection of arcs on $C$ whose both ends are defined by the vertices of $C$ such that no two arcs in $\mathcal{R}$ contain the same set of vertices.
@@
Let $\mathcal{R}$ be a collection of arcs on a cycle $C$ that are strict-containment free. Starting at an arbitrary vertex on $C$, in a clockwise traversal, we construct a cycle $C(\mathcal{R})$ on $\mathcal{R}$ as follows: When we visit a vertex $v$ of $C$, for each $R\in\mathcal{R}$ s.t. $s(R)=v$, we put a vertex in $C(\mathcal{R})$. The vertices added to $C(\mathcal{R})$ w.r.t. $v$ are ordered in increasing order of $t(R)$, with ties broken arbitrarily. We call the cycle $C(\mathcal{R})$ thus constructed, the cycle in lex. cyclic order, and we say that $\mathcal{R}$ is ordered in lex. cyclic order.
@@
$C(\mathcal{H})$ is clearly an outerplanar graph. Consider a subgraph $K\in\mathcal{K}$. Since $K$ induces a single run on $C$ and the subgraphs in $\mathcal{H}$ are strict-containment free, it follows that the subgraphs in $\mathcal{H}_{K}$ appear consecutively in lex. cyclic order on $C(\mathcal{H})$ and thus induce a connected subgraph of $C(\mathcal{H})$.
@@
The proof for the moreover part is similar. Indeed, let $h^{\prime}\in\mathrm{arc}[v,u]$ s.t. $h^{\prime}\in H^{\prime}\setminus H$. Also, note that $x\in H^{\prime}\setminus H$. Hence, the subgraphs $H,H^{\prime}$ form an $axax$-pattern as witnessed by the cyclic sequence $h^{\prime},u,x,v$; a contradiction.
@@
Note that since $\mathcal{H}$ is strict-containment free, so are $\mathcal{H}_{L}$ and $\mathcal{H}_{R}$, and that each $H\in\mathcal{H}\setminus\{H_{1},H^{\prime}_{s}\}$ contributes a run to exactly one of $\mathcal{H}_{L}$ or $\mathcal{H}_{R}$, by construction.
@@
((b)) $(C,\mathcal{H}_{L})$
((c)) $(C,\mathcal{H}_{R})$
Figure 5: Collections $\mathcal{H}_{L}$ and $\mathcal{H}_{R}$ of derived systems. $\mathcal{K}$ subgraphs are not shown in the figure.
@@
Finally, let $K$ intersect both $\mathrm{arc}[u_{0},v_{0}]$ and $\mathrm{arc}[v_{0},u_{0}]$. Note that in this case $K$ intersects $H_{0}$ (and hence with $H^{\prime}_{0}$ and $H^{\prime\prime}_{0}$) since $u_{0},v_{0}\in H_{0}$ and $(C,\mathcal{H},\mathcal{K})$ is strong $axax$-free. Each $H\in\mathcal{H}\setminus\{H_{0}\}$ has a representative vertex in one of $Q_{L}$ or $Q_{R}$ which are supports for the derived cycle systems with $H_{0}$ in both $Q_{L}$ and $Q_{R}$. It follows that $\mathcal{H}_{K}$ induces a connected subgraph of $Q$. Since $K$ was chosen arbitrarily, $Q$ is a support.
@@ 5.3 Implementation
Hence, the VC-dimension of the set system $(C,\mathcal{H})$ is at most 4, and this implies, by the Sauer-Shelah lemma (See Chapter 10 in [Mat02LecDiscGeom]) that $|\mathcal{H}|,|\mathcal{K}|=O(n^{4})$.
@@ 5.4 Primal and Dual Supports: Revisited
The existence of an intersection support also implies a primal and a dual support. However, we emphasize here that for a primal and a dual support, we require strictly weaker conditions. In particular, for a primal support, we require the cycle system $(C,\mathcal{H})$ to be $abab$-free and not necessarily $axax$-free, and for a dual support, we require $(C,\mathcal{H})$ to be $axax$-free. The result for a primal support directly follows from the following result of Raman and Singh [raman2025supportsgraphsboundedgenus].
@@
Theorem 25.
Let $(C,\mathcal{H})$ be an $abab$-free cycle system and $c:V(C)\to\{\mathbf{r},\mathbf{b}\}$ be a 2-coloring of the vertices of $C$. Then there is a primal support for $(C,\mathcal{H})$ that is an outerplanar graph.
Proof.
For each $v\in V(C)$ such that $c(v)=\mathbf{r}$, we delete it and make its neighbours adjacent. Let $C^{\prime}$ be the resulting cycle obtained form $C$, and =$\mathcal{H}^{\prime}=\{H\cap C^{\prime}:H\in\mathcal{H}\}$. Then $(C^{\prime},\mathcal{H}^{\prime})$ is an $abab$-free cycle system. By Lemma 24, there is a set $D$ of non-intersecting chord added to $C^{\prime}$ such that each subgraph in $\mathcal{H}^{\prime}$ induces a connected subgraph of $C^{\prime}\cup D$. Thus $C^{\prime}\cup D$ is an outerplanar primal support for $(C^{\prime},\mathcal{H}^{\prime})$. Note that this is also a primal support for $(C,\mathcal{H})$.
For the dual setting, one may likewise hope that the $abab$-free condition is sufficient to obtain a dual support. Unfortunately, that is not the case as the following example shows: Let $G$ be the graph on $\{1,2,\ldots 6\}$ as shown in Figure 6. Let $\mathcal{H}$ be subgraphs $H_{1}=\{1,2,3\}$, $H_{2}=\{3,4,5\}$, $H_{3}=\{5,6,1\}$ and $H_{4}=\{2,4,6\}$. It can be checked that $\mathcal{H}$ is $abab$-free. The dual support for $(G,\mathcal{H})$ is $K_{4}$, a complete graph on 4 vertices, which is not outerplanar.
The reason why $(G,\mathcal{H})$ does not admit a dual outerplanar support is that the induced system $(C,\mathcal{H})$ is not $axax$-free - $H_{4}$ forms an $axax$-pair with each of $H_{1},H_{2}$ and $H_{3}$. However, if $axax$-freeness is a sufficient condition for $(C,\mathcal{H})$ to exhibit an outerplanar dual support and it follows from Theorem 14.
Theorem 26.
If $(C,\mathcal{H})$ is $axax$-free, then it admits an outerplanar dual support.
Proof.
Let $\mathcal{K}=\{\{v\}:v\in V(C)\}$ be a collection of subgraphs each consisting of a single vertex. Then $(C,\mathcal{K})$ is $axax$-free. Therefore, the intersection system $(C,\mathcal{H},\mathcal{K})$ is $axax$-free. Note that the strong $axax$-free property is trivially satisfied by $(C,\mathcal{H},\mathcal{K})$. By Theorem 14, it has an intersection support for $Q$, which is also a dual support for $(C,\mathcal{H})$ since each $v\in V(C)$ corresponds to a subgraph in $K\in\mathcal{K}$.
@@ 6 Support of Bounded Treewidth
The proofs follow immediately from the fact that $\mathcal{H}$ are non-piercing and the adhesion set corresponding to any edge in $E(T)$ is a separator in $G$.
@@
Since $G$ is chordal, each bag induces a complete subgraph of $G$. Hence, each $H_{x}\in\mathcal{H}_{x}$ is connected. If $x$ is not a root of $T$, let $y$ be the parent of $x$. Then, for $H_{x},H^{\prime}_{x}\in\mathcal{H}_{x}$, the subgraph $H_{x}\setminus H^{\prime}_{x}$ of $G_{x}$ is connected since $H\setminus H^{\prime}$ is connected and $A_{xy}$ is a separator in $G$. Therefore, $(G_{x},\mathcal{H}_{x})$ is a non-piercing graph system. Finally, that $(T_{x},\mathcal{B}_{x})$ is a tree decomposition, is straightforward.
@@ 6.1 Primal Support
In this section, we show that a bounded treewidth non-piercing system $(G,\mathcal{H})$ with $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ admits a primal support $Q$ s.t. $\textsc{tw}(Q)\leq 2^{\textsc{tw}(G)+2}+\textsc{tw}(G)$. The proof is algorithmic and yields a polynomial time algorithm if $\textsc{tw}(G)$ is bounded, i.e., an FPT-algorithm parameterized by $\textsc{tw}(G)$.
@@
Let $(T,\mathcal{B})$ be a tree decomposition of $G$ of width $t$ where $T$ is a binary rooted tree. If $(T,\mathcal{B})$ is an easy tree decomposition, we are done by Lemma 29. Otherwise, we transform it into an easy tree decomposition of width at most $2^{t+2}+t$ by Lemma 30. We then obtain a primal support $Q$ of the same treewidth by Lemma 29.
@@ 6.2 Dual Support
We now construct a dual support. For a non-piercing system $(G,\mathcal{H})$ of treewidth $t$, we construct a dual support of treewidth $O(2^{4t})$. By Lemma 5, we assume that there are no containments, i.e., there are no two subgraphs $H,H^{\prime}\in\mathcal{H}$ such that $H\subseteq H^{\prime}$. As in the primal setting, we first show how we can obtain a dual support for a simple case. A support for the general case is obtained by reducing it to the simple case. In the following, since we construct a graph on $\mathcal{H}$, we abuse notation and also use $H$ to denote the vertex in the dual support corresponding to $H\in\mathcal{H}$.
@@
We now argue that $Q^{*}$ is a dual support for $(G,\mathcal{H})$. Consider a vertex $v\in V(G)$. The algorithm above ensures that each subgraph is pushed at most once. Suppose $H^{\prime}\in\mathcal{H}_{v}$ was pushed by $H$ so that its modified copy $H^{\prime\prime}=H^{\prime}\setminus H$ does not cover $v$. Then, $H\in\mathcal{H}_{v}$. Let $e=\{a,b\}$ be the connecting edge between $H$ and $H^{\prime\prime}$ such that $a\in H$ and $b\in H^{\prime\prime}$. Since $(T,\mathcal{B})$ is a valid tree-decomposition, there is a bag $B$ containing both $a$ and $b$. Let $H_{1}$ be the unique representative of $H^{\prime\prime}$ in $\textrm{unique}(\mathcal{H}^{\prime})$. Since Algorithm $k$-SDS puts a complete graph on the subgraphs intersecting $B$, it implies $H$ and $H_{1}$ are adjacent in $Q^{\prime}$. In $Q^{*}$, we made $H^{\prime}$ adjacent to $H_{1}$. Hence, $Q^{*}$ is a dual support for $(G,\mathcal{H})$.
@@
Let $(T,\mathcal{B})$ be a tree decomposition of $G$ of width $t$. If $(T,\mathcal{B})$ is $2^{4(t+1)}$-sparse, then we are done. Otherwise, by Lemma 32, we obtain a dual support $Q^{*}$ for $(G,\mathcal{H})$, of treewidth at most $2^{4(t+1)}$.